# VergeIO: Depth-Aware Eye Interaction on Glasses

### v1

@@ Abstract.
There is growing industry interest in creating unobtrusive designs for electrooculography (EOG) sensing of eye gestures on glasses (e.g. JINS MEME and Apple eyewear). We present VergeIO, the first EOG-based glasses that enables depth-aware eye interaction using vergence with an optimized electrode layout and novel smart glass prototype. It can distinguish between four and six depth-based eye gestures with 83–98% accuracy using personalized models in a user study across 11 users and 1,320 gesture instances. It generalizes to unseen users with an accuracy of 80–98% without any calibration. To reduce false detections, we incorporate a motion artifact detection pipeline and a preamble-based activation scheme. The system uses dry sensors without any adhesives or gel, and operates in real time with 3 mW power consumption by the sensing front-end, making it suitable for always-on sensing.
copyright: acmlicensedjournalyear: 2026submissionid: 13doi: XXXXXXX.XXXXXXXconference: The 24th ACM/IEEE Conference on Embedded Artificial Intelligence and Sensing Systems; May 6–9, 2026; Irvine, USAbooktitle: The 24th ACM/IEEE Conference on Embedded Artificial Intelligence and Sensing Systems (SenSys ’26), May 6–9, 2026, Irvine, USA
@@ 1. Introduction
Figure 1. VergeIO senses eye vergence on glasses using electrooculography (EOG) by detecting gaze shifts between objects at different depths. The eyes diverge when shifting focus from a near to a far object, and converge when focusing from a far to a near object.
Given their proximity to the eyes, glasses are uniquely positioned to leverage eye movements for interaction. A less-explored movement is vergence: the inward or outward motion of the eyes that occurs naturally as humans shift their gaze between objects at different depths (Fig. 1[1]. Unlike fast, reflexive saccades or lateral gaze shifts, vergence is slower and more deliberate, making it well-suited for intentional, hands-free interaction [2, 3, 4].
Figure 2. Applications of VergeIO. (a) Varifocal lenses adjust focal depth as the user shifts gaze between ambient reference points: a nearby phone and a distant TV. (b) Smart glasses enable gaze-based selection between virtual objects at different depths. (c) Remote screening of ocular disorders using approximate reference points: thumbs and a distant object. (d) Safety monitoring for drivers by detecting abrupt vergence shifts from the road to a nearby phone.
@@
  • Approximate physical depth cues. Remote screening for ocular disorders (e.g., convergence insufficiency, strabismus [5]) can be performed using one’s thumbs and a distant object as approximate depth markers.
Electrooculography (EOG) is a promising modality for vergence sensing on glasses, with growing industry interest in integrating EOG electrodes into glasses. Notably, these sensors have already been integrated into the nose pads and bridge of commercial-grade glasses including JINS MEME [6], and more recently, in late April 2025, Apple was granted a patent [7] to embed EOG sensors into glasses.
Unlike cameras, EOG is robust to varying ambient lighting conditions as glasses are worn across a range of environments. EOG is also low-power compared to commercial-grade cameras. In contrast to mmWave [8] and acoustic [9] systems that place sensors in front of the lenses and can obstruct the user’s vision, EOG electrodes have been integrated unobtrusively into the frames of the glasses.
The challenge, however, is that current EOG systems for glasses are not designed to robustly detect vergence. Specifically the electrodes on JINS MEME are placed too closely together to robustly detect vergence, which produces opposing electrical polarities from each eye and can cancel out or become indistinguishable [10]. As a result, such systems are typically limited to detecting blinks and coarse gaze shifts (e.g., left, right, up, down).
We present VergeIO, the first EOG-based glasses that expands the range of detectable eye movements to include distinct vergence gestures without compromising usability. Our key insight is that because EOG signals measure the eye’s electric field potential [11], reliably sensing the asymmetric movements of vergence requires a wider electrode separation. We leverage a two-channel electrode placement strategy that maximizes spatial separation, and present a novel hardware design that supports these placements, while remaining compatible with glasses.
@@
Importantly, vergence-based interaction can in principle extend across the spectrum of glasses, from high-end augmented reality smart glasses to basic audio-only frames like Amazon Echo Frames [12], or even instrumented varifocal glasses.
@@
  • We present a depth-aware glasses system which can distinguish between four and six unique vergence gestures with an accuracy ranging from 83 to 98% in an IRB-approved study across $n=11$ users using a personalized model. Our model generalizes to unseen users with an accuracy of 80 to 98% without additional calibration.
@@ 2. Background and related work
Reference Sensing Modality Vergence Sensing Robust to Ambient Light Unobstructed View Glasses Form Factor Power Consumption
CIDER [13] Cameras 40 µW
Tobii [14] Cameras DNS
FocusFlow [4] Cameras DNS
Li et al. [15] Photodiodes 395 µW
GazeTrak [9] Acoustic 16.4 mW
mmET [8] mmWave 6 mW
JINS MEME [6] EOG DNS
Mishra et al. [5] EOG DNS
@@
Eye sensing on headworn devices. There has been increased interest in eye movement sensing on head-mounted devices. Camera-based approaches [13, 16, 17, 14, 18, 19], while accurate, are power-hungry and are typically ineffective in the dark without infrared illumination. Photodiode-based systems [15], though lower in power, can be saturated by ambient sunlight in outdoor settings. Emerging acoustic [9] and mmWave systems [8] offer a promising low-power alternative, but can obstruct the user’s line of sight.
@@
Electrode-based approaches have made their way into VR headsets [20], smart glasses [21, 22, 23, 24, 25, 26, 27, 28], and earable devices [29, 30, 31]. They are often able to concurrently record electroencephalography (EEG), electrooculography (EOG), and electromyography (EMG) signals.
@@
Dry electrodes have been integrated into several commercial smart‐glasses platforms such as JINS MEME [6] and AttentivU [22] due to their small size, weight, and low power-consumption. They have achieved reliable blink detection and horizontal/vertical saccade tracking [10], facial‐expression recognition [24], midair gesture and context recognition [32], fatigue estimation [33], and mediated attention monitoring with real‐time biofeedback [34]. We build on this rich body of prior work by enabling vergence and depth-awareness as a key interaction primitive for EOG-based glasses.
@@
EOG-based systems. EOG has been used to detect various eye movements, including vergence, saccades, blinks, gaze shifts, and fixations [35, 36]. Prior systems have leveraged EOG-detected eye movements as an input modality for assistive interfaces for users with physical disabilities [37], robotic control [38], and hands-free operation of wearable computers for caregivers [39]. Beyond eye movement sensing, EOG has also been used to detect facial expressions (e.g., movements of the cheek, brow, and nose) [24] and to activity recognition such as typing, reading, eating, and talking [40].
@@
As EOG only requires small dry electrodes, there has been growing interest from both academia and industry to incorporate them into glasses [21, 27, 5]. Notable commercial efforts include JINS MEME [6], and Apple’s recently approved patent [7] to incorporate EOG electrodes into glasses. However currently proposed electrode configurations are either not able or not focused on sensing vergence, and are limited to blinks, winks, saccades and directional gaze shifts; or require a large number of electrodes that can be uncomfortable. In contrast, VergeIO expands the set of eye movements that can be detected on glasses to include six different vergence gestures using an electrode placement and a novel hardware glasses design that is optimized to capture the horizontal movements of the eye associated with vergence.
@@
Vergence-based systems. Vergence has emerged as a reliable eye-based input modality in extended reality environments [2, 4] due to its ability to mitigate the “Midas touch” problem, which refers to the difficulty of separating deliberate gaze commands from involuntary natural exploration as the eyes move around a scene [41]. Prior work has used eye-tracking cameras to measure vergence in VR systems for selecting and activating virtual objects [2, 4], controlling linear actuators [42], and studying visual attention in children with reading difficulties [43]. EOG-based systems for vergence sensing either rely on an electrode configuration that is incompatible with the form factor of glasses [44] or rely on customized flexible electrode designs [5] that specialized fabrication, making them more challenging to scale. In contrast, our approach focuses on detecting vergence using an EOG-based electrode configuration for glasses using conventional Ag/AgCl dry electrodes.
@@ 3. System design
Figure 3. Relationship between vergence distance and angle. (a) VergeIO measures changes in vergence angle ($\theta$) using EOG signals, which vary proportionally with these changes. (b) Our system can distinguish changes in vergence angles corresponding to three interaction distances: 30, 70, and 200 cm.
@@ 3.1. Vergence distance selection
A key system design parameter is the selection of vergence distances which is the physical distance between the eye and a fixation distance. These distances must satisfy two requirements: First, they should span the range of realistic, everyday viewing distances, and second, changes in vergence between these distances should result in discernible EOG signals that can be classified.
@@
Using an interpupillary distance (IPD) of 50 mm which is within the typical adult IPD range of 50-75 mm [45, 46], we select three vergence angles: 9.5°, 4.1°, and 1.4°. These yield uniformly spaced pairwise differences: 2.7°, 5.4°, and 8.1°, which enable them to be distinguishable in practice.
We show in Fig. 4 that the mean EOG curves for each of the six gestures when performed by using our proposed electrode placement by one participant. The figure shows that the curves are visually distinct. Additionally, we show in Fig. 3b that because the relationship between vergence angle and distance is non-linear, these vergence angles correspond to three vergence distances at key interaction distances: near (30 cm), mid (70 cm), and far (200 cm).
Figure 4. Mean horizontal EOG signal for each of the six vergence gestures for the left and right eye. Shaded region represents one standard deviation from the mean.
@@ 3.2. Placement of EOG electrodes
Electrode placement is a critical consideration in our design as it directly impacts the types of eye movements that can be detected. For instance, JINS MEME [6] is a commercial smart glass that integrates EOG sensing, but uses an electrode placement which cannot robustly sense vergence, as explicitly confirmed by prior work [10] authored by researchers from JINS MEME.
@@
To address these limitations, our system is designed to support vergence sensing while being compatible with a glasses form factor. Our design (Fig. 5b) places positive electrodes at each temple (left and right channels), a shared negative electrode on the nose bridge, and a ground electrode at the mastoid following best practices from prior work on EOG sensing [47]. By placing electrodes near the temples, our system expands the sensing area and is more sensitive to the asymmetric eye changes during vergence.
Figure 5. Limitations of prior EOG electrode placements on glasses. (a) JINS MEME’s design [6] lacks spatial separation which hinders vergence detection. Source: https://www.ohmyglasses.jp/blog/2016/05/31/jins-meme-new-world/ (b) VergeIO places electrodes at the temples and nose bridge which increases sensitivity of vergence detection. Flexible contact arms for the electrodes ensure stable skin contact across users while being compatible with a glasses form factor.
To empirically validate the advantage of our approach, we conduct a pilot study to measure the signal-to-noise ratio (SNR) of vergence gestures captured using the JINS MEME and VergeIO configuration. In this study, a single participant simultaneously wore electrodes placed according to both setups: specifically by wearing two electrodes on the temples, two on the nose pads, a shared electrode on the nose-bridge as the reference, and a ground electrode at the mastoid.
@@
Figure 7. SNR comparison across six vergence gestures, showing that VergeIO outperforms the commercial JINS MEME glasses [6].
We recorded four-channel EOG data over ten rounds, where each round consisted of six vergence gestures between the three distances we selected in Sec. 3.1. Synchronized data were sampled at 500 Hz and processed in sequence with a 3-order zero-phase low-pass filter at 10 Hz, a notch filter, and a third-order Savitzky–Golay filter (window size is 0.5 second). We manually segmented vergence events from the filtered signals of our two channels; for JINS MEME’s channels—where the gestures were nearly imperceptible—we used our channels’ onset–offset times to extract the corresponding segments. For each event, we extracted the two 0.2-second signals immediately before and after the event as noise segments. SNR was then computed as the ratio of the root-mean-square amplitude of the event segment to that of the concatenated noise segments.
We note that while prior work [48] on EOG sensing incorporates vertical electrodes above and below each eye, vergence is primarily a horizontal movement, and vertical channels only provide modest performance gains. Additionally, we report in our evaluation that (a) the addition of vertical electrodes were reported to induce discomfort by the participants (Sec. 5.1) and (b) our electrode configuration is still able to classify between different vertical eye movements (Sec. 5.4).

### v2

@@ Abstract.
There is growing industry interest in creating unobtrusive designs for electrooculography (EOG) sensing of eye gestures on glasses (e.g. JINS MEME and Apple eyewear). We present VergeIO, the first EOG-based glasses that enables depth-aware eye interaction using vergence with an optimized electrode layout and novel smart glass prototype. It can distinguish between four and six depth-based eye gestures with 83–98% accuracy using personalized models in a user study across 20 users and 2,400 gesture instances. It generalizes to unseen users with an accuracy of 77–97% without any calibration. To reduce false detections, we incorporate a motion artifact detection pipeline and a preamble-based activation scheme. The system uses dry sensors without any adhesives or gel, and operates in real time with 3 mW power consumption by the sensing front-end, making it suitable for always-on sensing.
vergence, eye sensing, electrooculography
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: IMWUT; Month DD–$D$, 2026; Shanghai, Chinaccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Figure 2. Applications of VergeIO. (a) Varifocal lenses adjust focal depth as the user shifts gaze between ambient reference points: a nearby phone and a distant TV. (b) Smart glasses enable gaze-based selection between virtual objects at different depths. (c) Remote screening of ocular disorders using approximate reference points: thumbs and a distant object.
Given their proximity to the eyes, glasses are uniquely positioned to leverage eye movements for interaction. A less-explored movement is vergence: the inward or outward motion of the eyes that occurs naturally as humans shift their gaze between objects at different depths (Fig. 1(Hughes, 1991). Unlike fast, reflexive saccades or lateral gaze shifts, vergence is slower and more deliberate, making it well-suited for intentional, hands-free interaction (Ahn et al., 2020; Kirst and Bulling, 2016; Zhang et al., 2024).
@@
  • Approximate physical depth cues. Remote screening for ocular disorders (e.g., convergence insufficiency, strabismus (Mishra et al., 2020)) can be performed using one’s thumbs and a distant object as approximate depth markers.
Electrooculography (EOG) is a promising modality for vergence (gaze depth) sensing on glasses, with growing industry interest in integrating EOG electrodes into glasses. Notably, these sensors have already been integrated into the nose pads and bridge of commercial-grade glasses including JINS MEME (jin, 2026), and more recently, Apple was granted a patent (pat, 2026) to embed EOG sensors into glasses.
Unlike cameras, EOG is robust to varying ambient lighting conditions as glasses are worn across a range of environments. EOG is also low-power compared to commercial-grade cameras. In contrast to mmWave (Ma et al., 2025) and acoustic (Li et al., 2024) systems that place sensors in front of the lenses and can obstruct the user’s vision, EOG electrodes have been integrated unobtrusively into the frames of the glasses. It is for these reasons that prior work (Schärer et al., 2025; Kosmyna et al., 2019; Frey et al., 2024; Lee et al., 2020; Kim et al., 2024; Rostaminia et al., 2019) has investigated EOG for capturing eye movements such as blinks and coarse gaze shifts (e.g., left, right, up, down) to support the understanding of interaction intentions.
The challenge, however, is that current EOG systems for glasses are not designed to robustly detect vergence. Specifically the electrodes on JINS MEME are placed too closely together to robustly detect vergence, which produces opposing electrical polarities from each eye and can cancel out or become indistinguishable (Uema and Inoue, 2017).
We present VergeIO, the first EOG-based glasses that expands the range of detectable eye movements to include distinct vergence gestures without compromising usability. Our key insight is that because EOG signals measure the eye’s electric field potential (Young and Sheena, 1975), reliably sensing the asymmetric movements of vergence requires a wider electrode separation. We leverage a two-channel electrode placement strategy that maximizes spatial separation, and present a novel hardware design that supports these placements, while remaining compatible with glasses.
@@
Importantly, vergence-based interaction can in principle extend across the spectrum of glasses, from high-end augmented reality smart glasses to basic audio-only frames like Amazon Echo Frames (ama, 2026), or even instrumented varifocal glasses.
@@
  • We present a depth-aware glasses system which can distinguish between four and six unique vergence gestures with an accuracy ranging from 83 to 98% in an IRB-approved study across $n=20$ users using a personalized model. Our model generalizes to unseen users with an accuracy of 77 to 97% without additional calibration. The system is compatible with traditional EOG sensing capability, achieving 96% accuracy across a 14-class set including blinks and coarse gaze shifts.
@@ 2. Background and related work
Reference Sensing Modality Vergence Sensing Robust to Ambient Light Unobstructed View Glasses Form Factor Power
CIDER (Mayberry et al., 2015) Cameras 40 µW
Tobii (tob, 2026) Cameras DNS
FocusFlow (Zhang et al., 2024) Cameras DNS
Li et al. (Li and Zhou, 2018) Photodiodes 395 µW
GazeTrak (Li et al., 2024) Acoustic 16.4 mW
mmET (Ma et al., 2025) mmWave 6 mW
JINS MEME (jin, 2026) EOG DNS
Mishra et al. (Mishra et al., 2020) EOG DNS
@@ 2.1. Eye movement sensing on headworn devices
There has been increased interest in eye movement sensing on head-mounted devices. Camera-based approaches (Mayberry et al., 2015, 2014; pup, 2026; tob, 2026; Cho et al., 2012; Ryan et al., 2008), while accurate, are power-hungry and are typically ineffective in the dark without infrared illumination. Photodiode-based systems (Li and Zhou, 2018), though lower in power, can be saturated by ambient sunlight in outdoor settings. Emerging acoustic (Li et al., 2024) and mmWave systems (Ma et al., 2025) offer a promising low-power alternative, but can obstruct the user’s line of sight.
@@
Electrode-based approaches have made their way into VR headsets (Pai et al., 2022), smart glasses (Frey et al., 2024; Kosmyna et al., 2019; Kanoh et al., 2015; Rostaminia et al., 2019; Kosmyna et al., 2024; Lee et al., 2020; Schärer et al., 2024; Kim et al., 2024), and earable devices (King et al., 2025; Liu et al., 2025; Lepold et al., 2024). They are often able to simultaneously record electroencephalography (EEG), electrooculography (EOG), and electromyography (EMG) signals.
@@
Dry electrodes have been integrated into several commercial smart‐glasses platforms such as JINS MEME (jin, 2026) and AttentivU (Kosmyna et al., 2019) due to their small size, weight, and low power-consumption. They have achieved reliable blink detection and horizontal/vertical saccade tracking (Uema and Inoue, 2017), facial‐expression recognition (Rostaminia et al., 2019), midair gesture and context recognition (Yeo et al., 2021), fatigue estimation (Tag et al., 2019), and mediated attention monitoring with real‐time biofeedback (Kosmyna et al., 2018). We build on this rich body of prior work by enabling vergence and depth-awareness as a key interaction primitive for EOG-based glasses.
2.2. EOG-based sensing systems
@@
EOG has been used to detect various eye movements, including vergence, saccades, blinks, gaze shifts, and fixations  (Belkhiria et al., 2022; Barea et al., 2002). However, it is important to note that EOG is generally not suitable for high-precision tracking, as its spatial resolution is typically limited to approximately 1.5° (Young and Sheena, 1975). Prior systems have leveraged EOG-detected eye movements as an input modality for assistive interfaces for users with physical disabilities (Patmore and Knapp, 1998), robotic control (Chen and Newman, 2004), and hands-free operation of wearable computers for caregivers (Mizuno et al., 2003). Beyond eye movement detection, EOG has also been used to detect facial expressions (e.g. movements of the cheek, brow, and nose) (Rostaminia et al., 2019) and to recognize activities such as typing, reading, eating, and talking (Ishimaru et al., 2014).
@@
As EOG only requires small dry electrodes, there has been growing interest from both academia and industry to incorporate them into glasses (Frey et al., 2024; Schärer et al., 2024; Mishra et al., 2020). Notable commercial efforts include JINS MEME (jin, 2026), and Apple’s recently approved patent (pat, 2026) to incorporate EOG electrodes into glasses. However, currently proposed electrode configurations are either not able or not focused on sensing vergence, and are limited to blinks, winks, saccades and directional gaze shifts; or require a large number of electrodes that can be uncomfortable. In contrast, VergeIO expands the set of eye movements that can be detected on glasses to include six different vergence gestures using an electrode placement and a novel hardware glasses design that is optimized to capture the horizontal movements of the eye associated with vergence.
2.3. Vergence as a interaction gesture
@@
Vergence has emerged as a reliable eye-based input modality in extended reality environments (Ahn et al., 2020; Zhang et al., 2024) due to its ability to mitigate the “Midas touch” problem, which refers to the difficulty of separating deliberate gaze commands from involuntary natural exploration as the eyes move around a scene (Jacob, 1990). Prior work has used eye-tracking cameras to measure vergence in VR systems for selecting and activating virtual objects (Ahn et al., 2020; Zhang et al., 2024), controlling linear actuators (Ruan et al., 2018), and studying visual attention in children with reading difficulties (Jiménez et al., 2020). EOG-based systems for vergence sensing either rely on an electrode configuration that is incompatible with the form factor of glasses (Tringali et al., 2021) or rely on customized flexible electrode designs (Mishra et al., 2020) that specialized fabrication, making them more challenging to scale. In contrast, our approach focuses on detecting vergence using an EOG-based electrode configuration for glasses using conventional Ag/AgCl dry electrodes.
@@ 3.1. Vergence distance selection
Given the ability of EOG to track eye movements with a precision of up to 1.5° (Young and Sheena, 1975), we define the depth-aware task for EOG as detecting transitions between discrete distances rather than performing continuous depth regression. Therefore, a key system design parameter is the selection of vergence distances which is the physical distance between the eye and a fixation distance. These distances must satisfy two requirements: First, they should span the range of realistic, everyday viewing distances, and second, changes in vergence between these distances should result in discernible EOG signals that can be classified.
Figure 3. Geometric characterization and signal validation for vergence distance selection. (a) VergeIO measures changes in vergence angle ($\theta$) using EOG signals, which vary proportionally with these changes. (b) Our system can distinguish changes in vergence angles corresponding to three interaction distances: 30, 70, and 200 cm. (c) Mean EOG signal for each of the six vergence gestures for the left and right eye. Shaded region represents one standard deviation from the mean.
@@
Using an interpupillary distance (IPD) of 50 mm which is within the typical adult IPD range of 50-75 mm (ipd, 2026; Dodgson, 2004), we select three vergence angles: 9.5°, 4.1°, and 1.4°. These yield uniformly spaced pairwise differences: 2.7°, 5.4°, and 8.1°, which enable them to be distinguishable in practice.
We show in Fig. 3c that the mean EOG curves for each of the six gestures when performed by using our proposed electrode placement by one participant. The figure shows that the curves are visually distinct. Additionally, we show in Fig. 3b that because the relationship between vergence angle and distance is non-linear, these vergence angles correspond to three vergence distances at key interaction distances: near (30 cm), mid (70 cm), and far (200 cm). The three distinct depth levels is motivated by both perceptual and practical considerations. Varifocal optical systems segment the visual field into near, intermediate, and far focal zones to accommodate the functional vision requirements recommended by the American Optometric Association (AOA) (Mancil et al., 2011). By matching this established discretization, the proposed systems naturally maps to the depth switching requirements of existing depth-based optical systems and typical viewing distances.
@@ 3.2. Placement of EOG electrodes
Electrode placement is a critical consideration in our design as it directly impacts the types of eye movements that can be detected. For instance, JINS MEME (jin, 2026) is a commercial smart glass that integrates EOG sensing, but uses an electrode placement which cannot robustly detect vergence, as explicitly confirmed by prior work (Uema and Inoue, 2017) authored by JINS MEME researchers.
@@
To address these limitations, our system is designed to support vergence sensing while being compatible with a glasses form factor. Our design (Fig. 4b) places positive electrodes at each temple (left and right channels), a shared negative electrode on the nose bridge, and a ground electrode at the mastoid following best practices from prior work on EOG sensing (Lopez et al., 2016). By placing electrodes near the temples, our system expands the sensing area and is more sensitive to the asymmetric eye changes during vergence. Additionally, by introducing a vertical separation between the temple and nose-bridge electrodes, our design also captures vertical eye movements. Consequently, VergeIO expands vergence sensing capability while maintaining compatibility with traditional horizontal and vertical EOG-based eye movements sensing, which was validated in Sec. 5.4.
Figure 4. Limitations of prior EOG electrode placements on glasses. (a) Design of JINS MEME (jin, 2026) and GAPses (Frey et al., 2024) lacks spatial separation which hinders vergence detection. Source: https://www.ohmyglasses.jp/blog/2016/05/31/jins-meme-new-world/ (b) VergeIO places electrodes at the temples and nose bridge which increases sensitivity of vergence detection. By introducing a vertical separation, the design is compatible with traditional horizontal and vertical EOG sensing. Flexible contact arms for the electrodes ensure stable skin contact across users while being compatible with a glasses form factor.
To empirically validate the advantage of our approach, we conduct a pilot study to measure the signal-to-noise ratio (SNR) of vergence gestures captured using the JINS MEME (jin, 2026) and GAPses (Frey et al., 2024) versus VergeIO configuration. In this study, a single participant simultaneously wore electrodes placed according to both setups: specifically by wearing two electrodes on the temples, two on the nose pads, a shared electrode on the nose-bridge as the reference, and a ground electrode at the mastoid.
Figure 5. SNR comparison across six vergence gestures. VergeIO outperforms the commercial JINS MEME glasses (jin, 2026).
We recorded four-channel EOG data over ten rounds, where each round consisted of six vergence gestures between the three distances we selected in Sec. 3.1. Synchronized data were sampled at 500 Hz and processed in sequence with a 3-order zero-phase low-pass filter at 10 Hz, a notch filter, and a third-order Savitzky–Golay filter (Savitzky and Golay, 1964) (window size is 0.5 second). We manually segmented vergence events from the filtered signals of our two channels; for JINS MEME’s channels—where the gestures were nearly imperceptible—we used our channels’ onset–offset times to extract the corresponding segments. For each event, we extracted the two 0.2-second signals immediately before and after the event as noise segments. SNR was then computed as the ratio of the root-mean-square amplitude of the event segment to that of the concatenated noise segments.
We note that while prior work (Narayan et al., 2024) on EOG sensing incorporates vertical electrodes above and below each eye, vergence is primarily a horizontal movement, and vertical channels only provide modest performance gains. Additionally, we report in our evaluation that (a) the addition of vertical electrodes were reported to induce discomfort by the participants (Sec. 5.1) and (b) our electrode configuration is still able to classify between different vertical eye movements (Sec. 5.4).